OpenStudy (rainbow_rocks03):

Will medal someone for best response so PLZ help. What is the area of a rectangle with a length of 7 4/53 3/4 feet?

OpenStudy (anonymous):

Multiply

OpenStudy (anonymous):

\(7 \dfrac{4}{5} \times 3 \dfrac{3}{4}\) First convert to improper fraction. 5 * 7 + 4 = ?

OpenStudy (rainbow_rocks03):

5*7+4=39

OpenStudy (rainbow_rocks03):

7 4/5=39 and 3 3/4=15/4

OpenStudy (rainbow_rocks03):

Help PLZ

OpenStudy (rainbow_rocks03):

HELP ME PLZ! someone